OpenStudy (firejay5):

Arithmetic Sequence: Find the indicated term of each arithmetic sequence. a of 1 = -4, d = -9, n = 20

OpenStudy (misty1212):

HI!!

OpenStudy (firejay5):

I got -175

OpenStudy (misty1212):

\[A-1=-4, d=-9, n=20\] so \(a_n=a_{20}=-4+19\times (-9)\)

OpenStudy (misty1212):

yeah \(-175\)

OpenStudy (firejay5):

thank you for checking for me